OpenStudy (anonymous):

Carlotta subscribes to the HotBurn music service. She can download no more than 11 song files per week. Carlotta has already downloaded 8 song files this week. Write, and solve an inequality to show how many more songs Carlotta can download.

OpenStudy (anonymous):

alright carlotta can only download 11 songs. lets dissect this sentence. We know she can have 11 song at the most and any thing less... so if we rewrite this, Carlotta can download less than or equal to 11. If we let this number be x we get \[x\leq11\]

OpenStudy (anonymous):

Now for the next sentence it says that she owns 8 songs already. This means that the number of songs she can download is the addition of her 8 songs + x(remember we allowed x = the amount of downloads . So rewriting this we get. 8+x must be less than or equal to 11 \[8+x\leq 11\]

OpenStudy (anonymous):

if you subtract 8 from both sides you can se \[x\leq 3\] which makes sense because if she's downloaded 8... she can download no more than 3 because if it's more than 3 say .... 4 8+4=12... which she's now getting charged =]